### v1

@@ III Diquarks and hidden charmed tetraquarks
To proceed the calculation, the constituent quark masses and the parameters in the $AL$ quark-quark potentials for the $cq$ (q=u, d) diquark and antidiquark are employed as those in Refs. [94, 95, 85].
In terms of those parameters, the masses of $cq$ diquarks and antidiquarks are computed with the $AL$ potentials, and the results are presented in Table 2. As a comparison, the corresponding masses from a QCD sum rule’s calculation [34] and a phenomenological analysis [30] are also presented.
Table 2: Mass (in MeV) of the $cq$ (q=u, d) diquark/antidiquark in $AL$ potentials.
@@
The $cq$ diquark with spin-$1$ has a mass ∼$\sim 43$ MeV larger than the $cq$ diquark with spin-00. The predicted mass of the $cq$ diquark with spin-$1$ is about ∼$\sim 300$ MeV larger than the corresponding one from QCD sum rule [34] or phenomenological analysis [30].
According to the argument in Ref. [85], the parameters $\alpha$ and $\lambda$ in the $AL$ quark-antiquark potential between a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark are fixed through two ground tetraquark candidates: $X(3872)$ and $T_{cc}(3875)^{+}$. $T_{cc}(3875)^{+}$ is assumed as a doubly charmed $J^{P}=1^{+}$ tetraquark made of a $cc$ diquark and an $\bar{u}\bar{d}$ antidiquark [85], and $X(3872)$ is assumed as a ground $J^{PC}=1^{++}$ hidden charmed tetraquark made from a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark. As indicated in Ref. [85], the real components of these two states(including mixing) will bring in a small uncertainty to the parameters $\alpha$ and $\lambda$. Four sets of $\alpha$ and $\lambda$ are also employed as those in Ref. [85]. The spin-spin interaction does not contribute to the masses of these two states, and the values of $\kappa$, $C$, $B$, $A$ and $r_{c}$ follow the original sets in Refs. [94, 95].
In terms of the parameters above and the obtained masses of diquarks/antidiquarks, the mass spectra of hidden charmed tetraquarks from $1S$ to $2P$ excitations are calculated. Masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-00 are presented in Table 3, masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$ are presented in Table 4, masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-$1$ are presented in Table 5.
Table 3: Mass spectrum (in MeV) of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-00 in $\mathbf{\bar{3}}_{c}\otimes\mathbf{3}_{c}$ color configuration.
@@
Table 4: Mass spectrum (in MeV) of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$ in $\mathbf{\bar{3}}_{c}\otimes\mathbf{3}_{c}$ color configuration.
@@
$1|[1,1]_{1},0\rangle_{1}$ $1^{+-}$ $3836.69$ $3835.79$ $3840.24$ $3839.88$ $4047$
$1|[1,1]_{2},0\rangle_{2}$ $2^{++}$ $3984.78$ $3986.44$ $3981.10$ $3982.54$
$1|[1,1]_{0},1\rangle_{1}$ $1^{--}$ $4250.16$ $4256.57$ $4248.70$ $4253.29$ $4545$
@@
[…]
$1|[1,1]_{1},1\rangle_{2}$ $2^{-+}$ $4284.59$ $4295.62$ $4287.76$ $4298.10$
$1|[1,1]_{2},1\rangle_{1}$ $1^{--}$ $4276.30$ $4281.85$ $4274.55$ $4279.97$ $4570$
$1|[1,1]_{2},1\rangle_{2}$ $2^{--}$ $4325.24$ $4334.72$ $4327.92$ $4336.65$
$1|[1,1]_{2},1\rangle_{3}$ $3^{--}$ $4333.34$ $4346.31$ $4338.96$ $4351.37$
@@
$2|[1,1]_{2},1\rangle_{2}$ $2^{--}$ $4744.90$ $4752.79$ $4753.13$ $4760.50$
$2|[1,1]_{2},1\rangle_{3}$ $3^{--}$ $4756.40$ $4767.28$ $4766.59$ $4776.99$
In these tables, the first column lists the tetraquark states in $LS$ coupling scheme from $1S$- to $2P$-wave excitations. The spin-parity quantum numbers $J^{PC}$ are given in the second column. The obtained mass spectra with the parameters sets I-IV of Ref. [85] are presented from the third to the sixth columns. The masses of hidden charmed tetraquarks obtained with four sets of parameters are very close.
As a comparison, the results obtained in the relativized diquark-antidiquark model [37] are also listed in the last column in these tables. The lowest hidden charmed tetraquark is the $0^{++}$ one with mass ∟$\sim 3832$ MeV, which is about $250$ MeV higher than the predicted one in Ref. [37], while masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$ are close to the predicted ones in that method. For the states consisting of a diquark and an antidiquark both with spin-$1$, our results cover a narrower mass range.
From Tables 3-5, we present explicitly some low excitations of hidden-charmed tetraquarks with different $J^{PC}$ in Table 6 for the convenience.
@@
$1^{++}$ $1|[1/0,0/1]_{1},0\rangle_{1}(\sim 3870)$ $2|[1/0,0/1]_{1},0\rangle_{1}(\sim 4430)$ X(3872)
$1^{--}$ $1|[0,0]_{0},1\rangle_{1}(\sim 4220)$ $1|[1/0,0/1]_{1},1\rangle_{1}(\sim 4250)$ $2|[0,0]_{0},1\rangle_{0}(\sim 4650)$ $2|[1/0,0/1]_{1},1\rangle_{1}(\sim 4680)$ $Y(4230)$, $X(Y)(4660)$
@@
In the table, the second to the fourth columns indicate the theoretical predictions, and the last column presents possible assignments of observed states. For tetraquarks without radial excitations ($n=1$), the lowest hidden charmed tetraquark has mass ∼$\sim 3770$ MeV with $J^{PC}=0^{++}$. This state consists of a diquark and an antidiquark both with spin-$1$. There is another $J^{PC}=0^{++}$ hidden-charmed tetraquark nearby with mass ∼$\sim 3830$ MeV, which has a diquark and an antidiquark both with spin-00. Their different inner structures may be distinguished from their hadronic decays. The mass of hidden-charmed tetraquark becomes higher in $1^{+-},~{}1^{++},~{}1^{--},~{}0^{-+},~{}0^{--},~{}1^{-+},~{}\cdots$ sequence. For tetraquarks with radial excitations ($n=2$), they have masses larger than $4300$ MeV.
@@
So far, many XYZ exotics have been observed. Some XYZ have been suggested as the normal $c\bar{c}$ charmoniums (see Table 1) though there are controversial opinions. Based on the $J^{PC}$ quantum numbers and masses (uncertainties considered) from Table 1, the following assignments of the observed XYZ exotics are natural suggestions.
@@
$X^{*}(3860)$ was first observed by Belle collaboration in the process $e^{+}e^{-}\to J/\Psi D\bar{D}$ [99], but not seen by LHCb collaboration in the decay $B^{+}\to D^{+}D^{-}K^{+}$ [100]. It could be assigned as the $0^{++}$ hidden charmed tetraquark composed of a diquark and an antidiquark both with spin-$1$. $Z_{c}(3900)$ and $X(3940)$ are possibly $1^{+-}$ hidden-charmed tetraquarks without radial excitation between the diquark and the antidiquark when theoretical uncertainties are taken into account. $X(3872)$ is fixed as the $1^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$. The $J$ and P quantum numbers of $X(4250)$ have not been determined. For its large mass uncertainty in experiment, $X(4250)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$ without radial excitation. If $X(4250)$ has mass around $4390$ MeV, it is possibly a $0^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark both with spin-00 with radial excitation. $X(4240)$ may be a $0^{--}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$.
@@
$X(4350)$ may be a $0^{++}$ hidden charmed tetraquark with radial excitation. $Z_{c}(4430)$ may be a $1^{+-}$ hidden-charmed tetraquark with radial excitation. The J and P quantum numbers of $X(4630)$ have not been determined yet, and there is also a large mass uncertainty for this state. $X(4630)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark with radial excitation. $X(Y)(4660)$ may be the $1^{--}$ hidden charmed tetraquarks with radial excitation.
@@
$Y(4260)$ was first observed by BaBar collaboration in the $\pi^{+}\pi^{-}J/\Psi$ mass spectrum [101], and late experiment for $e^{+}e^{-}\to\pi^{+}\pi^{-}J/\Psi$ by BESIII collaboration in a higher-statistics analysis [102] indicates that the $Y(4260)$ may consist of two resonances $Y(4230)$ and $Y(4320)$. The $Y(4230)$ is possibly a $1^{--}$ hidden charmed tetraquark without radial excitation. $Y(4390)$ was observed in the processes $e^{+}e^{-}\to\pi^{+}\pi^{-}\Psi(3686)$ by BESIII [103], whether it is the $Y(4360)$ is unknown yet. $Y(4008)$ was first observed by Belle Collaboration [104], but was not confirmed by subsequent experiments by BaBar and BESIII in the same production process. $Y(4008)$ was not thought as a genuine resonance, but a peak generated by the $\Psi(4040)$ with $D^{*}D$ loops with $\pi^{+}\pi^{-}J/\Psi$ in the final state [105]. From the numerical results, there exists no $1^{--}$ hidden charmed vector tetraquark located between $4300$ MeV and $4600$ MeV, or lower than $4200$ MeV. So $Y(4008)$ or $Y(4390)$ seems impossibly the $1^{--}$ hidden charmed tetraquark. Of course, more decay and production features are required for the final identification of these exotics.
@@
In order to find the spectra pattern of the hidden charmed tetraquarks, we have presented explicitly the mass splittings among different multiplets in Tables 7-9, respectively. The spin-weighted average mass of each multiplet is given by
@@
For the mass splittings among different charmonium in the last column in Tables 7-9, the masses of relevant charmonium are listed in Table 10, where the masses of well established $1S$-, $1P$- and $2S$-wave charmonia are taken from experimental data [8] while those of $2P$-wave charmonia are extracted from Ref. [106].
Table 7: Mass splittings (in MeV) of radially and orbitally excited $[cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark both with spin-00, where the parameters are chosen as Set I-IV.
@@
$1P-1S$ $387.21$ $396.24$ $389.29$ $397.51$ $456.7$
$2S-1S$ $552.53$ $557.68$ $561.82$ $567.06$ $605.4$
$2P-1P$ $428.06$ $426.62$ $433.03$ $431.86$ $415.4$
$2P-2S$ $262.74$ $265.18$ $260.50$ $262.31$ $266.8$
Table 8: Mass splittings (in MeV) of radially and orbitally $[cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$, where the parameters are chosen as Set I-IV.
@@
$1P-1S$ $382.07$ $385.47$ $384.90$ $390.02$ $456.7$
$2S-1S$ $551.60$ $556.83$ $561.82$ $567.06$ $605.4$
$2P-1P$ $425.49$ $421.99$ $432.21$ $430.10$ $415.4$
$2P-2S$ $255.96$ $250.63$ $255.29$ $253.06$ $266.8$
Table 9: Mass splittings (in MeV) of radially and orbitally $[cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark both with spin-$1$, where the parameters are chosen as Set I-IV.
@@
$2^{3}P_{2}$ $2^{++}$ $\chi_{c2}(2P)$ $3972$
As shown in Tables 7-9, the corresponding mass splittings of three kinds of hidden charmed tetraquarks are very close. The mass splittings of $1P-1S$ and $2S-1S$ are about $390-400$ MeV and $550-570$ MeV, respectively, which are about $70$ MeV and $50$ MeV smaller than those of charmonium. However, the mass splittings of $2P-1P$ and $2P-2S$ of hidden charmed tetraquarks are close to those of charmonium, which are about $250-260$ MeV and $420-430$ MeV, respectively. The feature for mass splittings exhibits a different spin-dependent contribution to the radial excitation in tetraquarks.
@@
In this work, a hidden charmed tetraquark without strange or antistrange quark is assumed consisting of a pair of $cq$ diquark and $\bar{c}\bar{q}$ antidiquark. The interactions in tetraquark are modulated as two kinds of interactions: the first type is the interaction between a quark and another quark, or the interaction between an antiquark and another antiquark; the second type is the interaction between a diquark and an antiquark. For the practical calculation, the modified $AL$-type quark-quark and quark-antiquark potentials proposed by Semay and Silvestre-Brac are employed. To describe the interaction within a tetraquark in a better way, the parameters of the potential between the diquark and the antidiquark are refitted with the experimental mass of $X(3872)$ and $T_{cc}(3875)^{+}$.
In terms of the Semay-Silvestre-Brac potentials, the mass of the $cq$ diquark or $\bar{c}\bar{q}$ antidiquark with spin-00 is predicted with ∼$\sim 2175$ MeV, and the mass of that with spin-$1$ is predicted with ∼$\sim 2220$ MeV. The mass splitting between the scalar and vector diquark is ∼$\sim 45$ MeV. The predicted mass of $cq$ diquark with spin-$1$ is about ∼$\sim 300$ MeV larger than the predicted one from QCD sum rule or phenomenological analysis.
Semay-Silvestre-Brac potentials are also employed to describe the interactions between the diquark and the antidiquark. When the parameters $\alpha$ and $\lambda$ are fixed by the two ground tetraquark candidates $X(3872)$ and $T_{cc}(3875)$, the masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are calculated. The lowest hidden charmed tetraquark is predicted with mass ∼$\sim 3770$ MeV and $J^{PC}=0^{++}$. The mass of hidden charmed tetraquark becomes higher in $1^{+-},~{}1^{++},~{}1^{--},~{}0^{-+},~{}0^{--},~{}1^{-+},~{}\cdots$ sequence, where the tetraquarks with exotic $J^{PC}=0^{--},~{}1^{-+}$ have higher masses. The hidden charmed tetraquark with radial excitation has mass larger than $4300$ MeV.
Based on numerical results of the tetraquarks, some observed XYZ exotics have been analysed and their possible $J^{PC}$ and inner structures are suggested. $X^{*}(3860)$ could be assigned as the $0^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark both with spin-$1$. $Z_{c}(3900)$ and $X(3940)$ are possibly the $1^{+-}$ hidden charmed tetraquarks without radial excitation. $X(3872)$ is fixed as a $1^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$. $X(4250)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark without radial excitation, or a $0^{++}$ hidden charmed tetraquark with radial excitation. $X(4240)$ may be a $0^{--}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-00 and one spin-$1$. $Y(4008)$ or $Y(4390)$ seems impossibly the $1^{--}$ hidden-charmed tetraquark.

### v2

@@ III Diquarks and hidden charmed tetraquarks
To proceed to the calculation, the constituent quark masses and the parameters in the $AL$ quark-quark potentials for the $cq$ (q=u, d) diquark and antidiquark are employed as those in Refs. [95, 96, 81].
In terms of those parameters in Refs. [95, 96], the masses of $cq$ diquarks and antidiquarks are computed with the $AL$ potentials, and the results are presented in Table 2. As a comparison, the corresponding masses from a phenomenological analysis [31] and a QCD sum rule calculation [35] are also presented.
Table 2: Mass (in MeV) of the $cq$ (q=u, d) diquark/antidiquark in ALAL potentials.
@@
The $cq$ diquark with spin-$1$ has a mass $\sim 43$ MeV greater than that of the $cq$ diquark with spin-0. The predicted mass of the $cq$ diquark with spin-$1$ is about $\sim 300$ MeV greater than the corresponding one from the phenomenological analysis [31] or the QCD sum rule [35].
In practical calculations of spectra of hidden charmed tetraquark states, the two parameters $\alpha$ and $\lambda$ employed in this work are slightly different from those for conventional mesons and baryons. According to the argument in Ref. [81], the parameters $\alpha$ and $\lambda$ in the $AL$ quark-antiquark potential between a $cq$ diquark and a $\bar{c}\bar{q}$ antidiquark are fixed through two ground state tetraquark candidates: $X(3872)$ and $T_{cc}(3875)^{+}$. $T_{cc}(3875)^{+}$ is assumed as a doubly charmed $J^{P}=1^{+}$ tetraquark made of a $cc$ diquark and a $\bar{u}\bar{d}$ antidiquark [81], and $X(3872)$ is assumed as a hidden charmed $J^{PC}=1^{++}$ ground state tetraquark made from a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark. As indicated in Ref. [81], the real components of these two states (including mixing) will bring in a small uncertainty to the parameters $\alpha$ and $\lambda$. A combination of two types of AL potentials applied to diquarks/antidiquarks leads to four sets of parameters. The four sets of $\alpha$ and $\lambda$ are employed as those in Ref. [81]. The spin-spin interaction does not contribute to the masses of these two states, and the values of $\kappa$, $C$, $B$, AA and $r_{c}$ follow the original sets in Refs. [95, 96].
In terms of the parameters shown above and the obtained masses of diquarks/antidiquarks, the mass spectra of hidden charmed tetraquarks from $1S$ to $2P$ excitations are calculated through the solution of matrix eigenvalue problem as done in Refs. [99, 81]. Masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-0 are presented in Table 3, masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$ are presented in Table 4, and masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-$1$ are presented in Table 5.
Table 3: Mass spectrum (in MeV) of hidden charmed tetraquarks consisting of a diquark and an antidiquark both with spin-0 in $\mathbf{\bar{3}}_{c}\otimes\mathbf{3}_{c}$ color configuration.
@@
Table 4: Mass spectrum (in MeV) of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$ in $\mathbf{\bar{3}}_{c}\otimes\mathbf{3}_{c}$ color configuration.
@@
$|[1,1]_{1},0\rangle_{1} 1+1^{+-}$ $3836.69$ $3835.79$ $3840.24$ $3839.88$ $4047$
$1|[1,1]_{2},0\rangle_{2}$ $2^{++}$ $3984.78$ $3986.44$ $3981.10$ $3982.54$ -
$|[1,1]_{0},1\rangle_{1} 11^{--}$ $4250.16$ $4256.57$ $4248.70$ $4253.29$ $4545$
@@
$|[1,1]_{1},1\rangle_{1} 1+1^{-+}$ $4279.19$ $4287.89$ $4280.40$ $4288.28$ -
$1|[1,1]_{1},1\rangle_{2}$ $2^{-+}$ $4284.59$ $4295.62$ $4287.76$ $4298.10$ -
$|[1,1]_{2},1\rangle_{1} 11^{--}$ $4276.30$ $4281.85$ $4274.55$ $4279.97$ $4570$
$1|[1,1]_{2},1\rangle_{2}$ $2^{--}$ $4325.24$ $4334.72$ $4327.92$ $4336.65$ -
$1|[1,1]_{2},1\rangle_{3}$ $3^{--}$ $4333.34$ $4346.31$ $4338.96$ $4351.37$ -
@@
$|[1,1]_{2},1\rangle_{2} 22^{--}$ $4744.90$ $4752.79$ $4753.13$ $4760.50$ -
$2|[1,1]_{2},1\rangle_{3}$ $3^{--}$ $4756.40$ $4767.28$ $4766.59$ $4776.99$ -
In these tables, the first column lists the tetraquark states in $LS$ coupling scheme from $1S$- to $2P$-wave excitations. The spin-parity quantum numbers $J^{PC}$ are given in the second column. The obtained mass spectra with the parameters sets I-IV of Ref. [81] are presented in the third to sixth columns. The masses of hidden charmed tetraquarks obtained with four sets of parameters are very close.
As a comparison, the results obtained in the relativized diquark-antidiquark model [38] are also listed in the last column in these tables. The lowest hidden charmed tetraquark is the $0^{++}$ one with mass $\sim 3832$ MeV, which is about $250$ MeV higher than the predicted one in Ref. [38], while masses of hidden charmed tetraquarks consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$ are close to the predicted ones in that method. For the states consisting of a diquark and an antidiquark both with spin-$1$, our results cover a narrower mass range.
From Tables 3-5, we present explicitly some low excitations of hidden charmed tetraquarks with different $J^{PC}$ in Table 6 for later convenience.
@@
$1^{++}$ $1|[1/0,0/1]_{1},0\rangle_{1}(\sim 3870)$ - $2|[1/0,0/1]_{1},0\rangle_{1}(\sim 4430)$ - X(3872)
$1^{--}$ $1|[0,0]_{0},1\rangle_{1}(\sim 4220)$ $1|[1/0,0/1]_{1},1\rangle_{1}(\sim 4250)$ $2|[0,0]_{0},1\rangle_{0}(\sim 4650)$ $2|[1/0,0/1]_{1},1\rangle_{1}(\sim 4680)$ Y(4230)Y(4230)$X(Y)(4660)$
@@
In Table 6, the second to the fourth columns indicate the theoretical predictions, and the last column presents possible assignments of observed states. For tetraquarks without radial excitations ($n=1$), the lowest hidden charmed tetraquark has mass $\sim 3770$ MeV with $J^{PC}=0^{++}$. This state consists of a diquark and an antidiquark both with spin-$1$. There is another hidden charmed $J^{PC}=0^{++}$ tetraquark nearby with mass $\sim 3830$ MeV, which has a diquark and an antidiquark both with spin-0. Their different inner structures may be distinguished from their hadronic decays. The mass of hidden-charmed tetraquark becomes higher in the $1^{+-},~1^{++},~1^{--},~0^{-+},~0^{--},~1^{-+}$, … sequence. For tetraquarks with radial excitations ($n=2$), they have masses greater than $4300$ MeV.
@@
So far, many XYZ exotics have been observed. Some XYZ exotics have been suggested as the normal $c\bar{c}$ charmoniums (see Table 1) though there are controversial opinions. Based on the quantum numbers $J^{PC}$ and masses (uncertainties considered) from Table 1, the following assignments of the observed XYZ exotics are natural suggestions.
@@
$X^{*}(3860)$ was first observed by the Belle collaboration in the process $e^{+}e^{-}\to J/\Psi D\bar{D}$ [100], but not seen by the LHCb collaboration in the decay $B^{+}\to D^{+}D^{-}K^{+}$ [101]. It could be assigned as the hidden charmed $0^{++}$ tetraquark composed of a diquark and an antidiquark both with spin-$1$. $Z_{c}(3900)$ and $X(3940)$ are possibly $1^{+-}$ hidden-charmed tetraquarks without radial excitation between the diquark and the antidiquark when theoretical uncertainties are taken into account. $X(3872)$ is fixed as the $1^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$. The quantum numbers $J$ and $P$ of $X(4250)$ have not been determined. For its large mass uncertainty in experiment, $X(4250)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$ without radial excitation. If $X(4250)$ has mass around $4390$ MeV, it is possibly a hidden charmed $0^{++}$ tetraquark consisting of a diquark and an antidiquark both with spin-0 with radial excitation. $X(4240)$ may be a $0^{--}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$.
@@
$X(4350)$ may be a $0^{++}$ hidden charmed tetraquark with radial excitation. $Z_{c}(4430)$ may be a hidden charmed $1^{+-}$ tetraquark with radial excitation. The quantum numbers $J$ and $P$ of $X(4630)$ have not been determined yet, and there is also a large mass uncertainty for this state. $X(4630)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark with radial excitation. $X(Y)(4660)$ may be the $1^{--}$ hidden charmed tetraquarks with radial excitation.
@@
$Y(4260)$ was first observed by the BaBar collaboration in the $\pi^{+}\pi^{-}J/\Psi$ mass spectrum [102], and late experiment for $e^{+}e^{-}\to\pi^{+}\pi^{-}J/\Psi$ by the BESIII collaboration in a higher-statistics analysis [103] indicates that $Y(4260)$ may consist of two resonances $Y(4230)$ and $Y(4320)$. $Y(4230)$ is possibly a $1^{--}$ hidden charmed tetraquark without radial excitation. $Y(4390)$ was observed in the processes $e^{+}e^{-}\to\pi^{+}\pi^{-}\Psi(3686)$ by the BESIII collaboration [104], and whether it is the $Y(4360)$ is unknown. $Y(4008)$ was first observed by the Belle Collaboration [105], but was not confirmed by subsequent experiments by the BaBar and BESIII collaborations in the same production process. $Y(4008)$ was not thought as a genuine resonance, but a peak generated by the $\Psi(4040)$ with $D^{*}D$ loops with $\pi^{+}\pi^{-}J/\Psi$ in the final state [106]. From the numerical results, there exists no $1^{--}$ hidden charmed vector tetraquark located between $4300$ and $4600$ MeV, or lower than $4200$ MeV. So it seems impossible for $Y(4008)$ or $Y(4390)$ to be the $1^{--}$ hidden charmed tetraquark. Of course, more decay and production features are required for the final identification of these exotics.
@@
In order to find the spectra pattern of the hidden charmed tetraquarks, we have presented explicitly the mass splittings among different multiplets in Tables 7-9. The spin-weighted average mass of each multiplet is given by
@@
For the mass splittings among different charmoniums in the last column in Tables 7-9, the masses of relevant charmoniums are listed in Table 10, where the masses of well established $1S$-, $1P$- and $2S$-wave charmoniums are taken from experimental data [8] while those of $2P$-wave charmoniums are extracted from Ref. [107].
Table 7: Mass splittings (in MeV) of radially and orbitally excited $cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark both with spin-0, where the parameters are chosen as Sets I-IV.
@@
1P1S1P-1S $387.21$ $396.24$ $389.65$ $397.80$ $456.7$
2S1S2S-1S $552.53$ $557.68$ $562.79$ $567.94$ $605.4$
2P1P2P-1P $428.06$ $426.62$ $434.15$ $432.94$ $415.4$
2P2S2P-2S $262.74$ $265.18$ $261.01$ $262.80$ $266.8$
Table 8: Mass splittings (in MeV) of radially and orbitally $cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark with one spin-0 and one spin-11, where the parameters are chosen as Sets I-IV.
@@
1P1S1P-1S $386.96$ $396.08$ $389.29$ $397.51$ $456.7$
2S1S2S-1S $551.60$ $556.83$ $561.82$ $567.06$ $605.4$
2P1P2P-1P $426.93$ $425.52$ $433.03$ $431.86$ $415.4$
2P2S2P-2S $262.29$ $264.77$ $260.50$ $262.31$ $266.8$
Table 9: Mass splittings (in MeV) of radially and orbitally $cq][\bar{c}\bar{q}]$ consisting of a diquark and an antidiquark both with spin-$1$, where the parameters are chosen as Sets I-IV.
@@
$^{3}P_{2} 2++2^{++}$ $\chi_{c2}(2P)$ $3972$
As shown in Tables 7-9, the corresponding mass splittings of three kinds of hidden charmed tetraquarks are very close. The mass splittings of 1P1S1P-1S and 2S1S2S-1S are about 390400390-400 and 550570550-570 MeV, respectively, which are about $50$ and $40$ MeV smaller than those of charmoniums. However, the mass splittings of 2P1P2P-1P and 2P2S2P-2S of hidden charmed tetraquarks are close to those of charmonium, which are about 250260250-260 MeV and 420430420-430 MeV, respectively. The feature for mass splittings exhibits a different spin-dependent contribution to the radial excitation in tetraquarks.
@@
In this work, a hidden charmed tetraquarks without strange or antistrange quark is assumed consisting of a pair of $cq$ diquark and $\bar{c}\bar{q}$ antidiquark. The interactions in tetraquark are modulated as two kinds of interactions: the first type is the interaction between a quark and another quark, or the interaction between an antiquark and another antiquark; the second type is the interaction between a diquark and an antiquark. For the practical calculation, the modified $AL$-type quark-quark and quark-antiquark potentials proposed by Semay and Silvestre-Brac are employed. To describe the interaction within a tetraquark in a better way, the parameters of the potential between the diquark and the antidiquark are refitted with the experimental mass of $X(3872)$ and $T_{cc}(3875)^{+}$.
In terms of the Semay-Silvestre-Brac potentials, the mass of the $cq$ diquark or $\bar{c}\bar{q}$ antidiquark with spin-0 is predicted with $\sim 2175$ MeV, and the mass of that with spin-$1$ is predicted with $\sim 2220$ MeV. The mass splitting between the scalar and vector diquark is $\sim 45$ MeV. The predicted mass of the $cq$ diquark with spin-$1$ is about $\sim 300$ MeV greater than the predicted one from QCD sum rule or phenomenological analysis.
Semay-Silvestre-Brac potentials are also employed to describe the interactions between the diquark and the antidiquark. When the parameters $\alpha$ and $\lambda$ are fixed by the two ground state tetraquark candidates $X(3872)$ and $T_{cc}(3875)$, the masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are calculated. The lowest hidden charmed tetraquark is predicted with mass $\sim 3770$ MeV and $J^{PC}=0^{++}$. The mass of hidden charmed tetraquark becomes higher in the $1^{+-},~1^{++},~1^{--},~0^{-+},~0^{--},~1^{-+}$, … sequence, where the tetraquarks with exotic $J^{PC}=0^{--},~1^{-+}$ have greater masses. The hidden charmed tetraquark with radial excitation has mass greater than $4300$ MeV.
Based on the numerical results of the tetraquarks, some observed XYZ exotics have been analyzed and their possible $J^{PC}$ and inner structures are suggested. $X^{*}(3860)$ could be assigned as the $0^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark both with spin-$1$. $Z_{c}(3900)$ and $X(3940)$ are possibly the $1^{+-}$ hidden charmed tetraquarks without radial excitation. $X(3872)$ is fixed as a $1^{++}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$. $X(4250)$ may be a $0^{-+}$ or $1^{-+}$ hidden charmed tetraquark without radial excitation, or a $0^{++}$ hidden charmed tetraquark with radial excitation. $X(4240)$ may be a $0^{--}$ hidden charmed tetraquark consisting of a diquark and an antidiquark with one spin-0 and one spin-$1$. $Y(4008)$ or $Y(4390)$ seems impossible to be the $1^{--}$ hidden charmed tetraquark.